OpenStudy (anonymous):

he equation 3x^2+2y^2=0 represents a a. circle b. parabola c. pair of lines d.none of the above

OpenStudy (anonymous):

circle

OpenStudy (anonymous):

but i think that the coefficients of x and y must be 1 when its a circle

OpenStudy (anonymous):

no actually it might be just one line. Try putting x on the other side and try to get y by itself. Then it looks more like a line

OpenStudy (andriod09):

the question is this correct? \[3x^{2}+2y^{2}=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ellipse with radius zero?

OpenStudy (andriod09):

lets see, on a graph when you have the coefficients x and y, they what?

OpenStudy (anonymous):

if there is no radius then it isn't an elipse

OpenStudy (anonymous):

That's my point.

OpenStudy (andriod09):

 they intersect, od then it can be a circle.

OpenStudy (anonymous):

then is it a cirlcle?

OpenStudy (andriod09):

can't*

OpenStudy (andriod09):

They're squared, so then it COULD be a parabula,

OpenStudy (andriod09):

OR they COULD be a pair on lines.

OpenStudy (andriod09):

i believe you have to solve it to find out. :)

OpenStudy (anonymous):

Solving for y: \[y= \pm \sqrt{\frac{-3x^2}{2}}\]

OpenStudy (anonymous):

It's a graph of the origin.

OpenStudy (andriod09):

you dont have to un sqaire it. you know. you just multiply it out. \[3x^{2}+2y^{2}=0\]

OpenStudy (anonymous):

Multiply what? It's already in simplest form... ("That's my point." <-- get this joke yet?)

OpenStudy (andriod09):

(Solve for x/y then substitute)

OpenStudy (anonymous):

Substitution would be redundant, except to verify the solution. Looking at the formula for y, it is clear that there is only one solution (the ellipse with radius = 0).

OpenStudy (andriod09):

i already know the answer, im helping him find it, an no its not. its a quadratic equation. so how can it be the ellipse with the radius=0?

OpenStudy (anonymous):

There is no algebra necessary here, only reasoning.

OpenStudy (anonymous):

1.std eqn. for circle is x^2+y^2=a^2 2.if h^2-ab=or>0 then it represents a pol thats why im confused

OpenStudy (anonymous):

What would the graph of an ellipse with radius = 0 look like?

OpenStudy (anonymous):

The coefficients 3 and 2 are irrelevant; it is two square numbers adding up to zero - what do those two numbers have to be then?

OpenStudy (andriod09):

ohh, but Cliff, this requires algebra, :) ik my math.

OpenStudy (anonymous):

Yes, you could use algebra to rearrange the equation like I did when I solved for y if that helps you see the solution easier, but it's not necessary. Experience with general quadratic equations and conic sections provide the standard form of an equation for an ellipse: ax^2 + by^2 = c where a, b, and c are non-negative. c affects the size of the ellipse. If c=0 then the ellipse has no size.

OpenStudy (anonymous):

y=sqrt(-1.5x^2) This is a line

OpenStudy (anonymous):

pair?

OpenStudy (anonymous):

@edwardo24 what?

OpenStudy (andriod09):

@edwardo24 you mean: \[y=\sqrt{-1.5x^{2} }\]

OpenStudy (anonymous):

yes that is what i mean

OpenStudy (anonymous):

(That's also the exact same equation I provided earlier, except you forgot the plus/minus)

OpenStudy (anonymous):

And that is *not* the equation for a line (unless you want to call it a line of zero length).

OpenStudy (anonymous):

its not a line of 0 length

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

If it is a line, then you should be able to put the equation in the form y=mx+b and tell me what the slope is.

OpenStudy (andriod09):

no, thats not the one you gave earlyer. the one you gave was this: \[y=\pm\sqrt{\frac{-3x^{2} } { 2 }}\]

OpenStudy (anonymous):

Has anyone even tried putting values of x into that function to see what happens?

OpenStudy (andriod09):

we don't know what x is, remember?

OpenStudy (anonymous):

I don't see why this is so difficult to understand. It is an ellipse of zero size. The graph of that equation is the origin.

OpenStudy (andriod09):

no its not.

OpenStudy (anonymous):

It's a single equation with two variables, that makes one of them a free parameter and you can choose values arbitrarily.

OpenStudy (anonymous):

Are you kidding me? Then tell me what it looks like then.

OpenStudy (andriod09):

look, this is annoying me, its a quadratic equation: \[x^{2}+y^{2}=0\] just with coefficients 3 and 2.

OpenStudy (anonymous):

Go on..

OpenStudy (andriod09):

it the perabula. not a freaken circle. if there are both \[x^{2}\] and \[y{2}\] that equal 0 in this manner, \[x^{2}+y^{2}=0\] then its the quadratic equation. just moved up 3 and to the left 2.

OpenStudy (anonymous):

Here's an x-y table so you can plot the solutions: